Maintaining a healthy weight is important for women’s health. Some physicians recommend being back to your pre-pregnancy weight six months after giving birth. Public health officials are interested in whether this is a realistic goal. Do the women in North Carolina weigh more than their pre-pregnancy weight six months postpartum? In other words, is the six months post-pregnancy weight greater than their pre-pregnancy weight? Use a paired samples t-test to compare pst6wght (variable 1) and prewght (variable 2). This uses pst6wght-prewght. (12 points)

1. Set-up the hypotheses and significance level (2 points) 2. Set-up the decision rule (1 point) 3. Compute the test statistic using Excel (the test statistic is in your Excel output.) (5 points) 4. Give your conclusion. This must include either a comparison of the p-value to alpha or a comparison of the test statistic to the critical value. You also must explain what this result means. (4 points)

PREWGHT PST6WGHT
180

195

204 195
135 140
107 130
144 144
120 125
139 135
187 190
158 185
200 204
127 124
134 132
210 201
110 115
192 191
163 165
148 149
105 115
217 218
182 195
111 115
222 203
159 161
217 203
157 149
123 125
206 201
116 119
206 193
220 201
148 145
123 124
160 167
203 201
168 154
117 123
197 186
191 195
223 225
166 189
171 201
223 225
111 141
105 107
150 152
107 137
114 129
146 148
150 149
217 204
207 220
198 192
99 124
131 166
125 142
169 168
163 164
154 157
128 143
174 191
219 198
100 106
118 124
184 175
113 113
181 185
136 145
187 184
180 181
189 191
205 205
224 201
215 216
111 119
124 121
156 186
124 128
199 195
158 165
155 158
153 152
214 215
173 185
209 211
162 165
176 165
159 162
166 154
118 118
149 157
151 196
101 102
182 189
155 163
119 121
107 108
196 194
203 201
115 128
144 149
126 154
115 121
147 156
205 210
160 160
225 252
139 137
153 139
184 181
182 185
111 124
182 191
144 145
127 127
170 174
123 136
119 125
221 229
195 201
211 242
204 235
98 105
136 135
144 142
160 168
183 181
156 159
99 125
126 130
107 117
225 190
221 210
123 136
140 146
140 141
134 137
123 141
222 204
222 221
194 195
171 172
108 125
113 124
103 118
98 113
115 131
134 154
167 171
131 135
214 249
188 187
212 213
203 205
113 133
203 201
101 104
182 181
119 121
99 119
176 210
134 136
198 195
137 138
105 108
147 156
146 147
219 215
176 208
114 119
209 210
125 158
147 148
177 178
168 186
115 116
101 111
147 157
113 121
112 125
179 185
124 125
120 126
185 195
126 146
136 145
206 208
212 204
124 145
148 154
101 106
207 204
185 199
121 143
146 152
163 178
208 203
137 145
136 139
153 175
175 165
Answer #1


Answer Date: 11/11/2019 To test the hypothesis is that the six months post-pregnancy weight is greater than their pre-pregnand.f.=n-1 = 200 -1 = 199 The t critical value at 5% significance level is 1.653 . The decision rule is, Reject the null hypothChoose Output Range from Output options. Click Ok. 7) t-Test: Paired Two Sample for Means PSTOWGHT 162.825 1242.868719 200 0.4) Decision The statistical decision is that the t-test statistics which is 6.407 is higher than t-critical value which is 1.